what is the solution to the system below?
y=5x-9
y=x+3

Answer by ikleyn(52788)

.

Since left sides are identical, right sides are equal


    5x - 9 = x + 3


Simplify and find x


    5x - x = 3 + 9

    4x     = 12

     x     = 12/4 = 3.


Then from either of the two given equations,  y = 6.


ANSWER.  The solution  is  (x,y) = (3,6).